# RMAAT: Astrocyte-Inspired Memory Compression and Replay for Efficient Long-Context Transformers

### v1

@@ 3.4 AMRB Training Algorithm
Training recurrent architectures on long sequences via standard BPTT is often memory-prohibitive due to storing activations for the entire sequence length. To overcome this while leveraging RMAAT’s unique memory structure, we introduce the Astrocytic Memory Replay Backpropagation (AMRB) algorithm (Contribution $3$), an efficient training approach inspired by techniques for recurrent networks (Bellec et al., 2019; Meng et al., 2023)The core idea of AMRB is to avoid storing all intermediate activations within each segment during the forward pass. Instead, it leverages the persistent, compressed Memory Tokens ($mem_{t}$) described in Section 3.3. During the forward pass through $T_{seg}$ segments, only the sequence of memory token states ($mem_{1},mem_{2},...,mem_{T_{seg}+1}$) passed between segments is stored in a replay buffer. During the backward pass, gradients are computed segment by segment. To calculate the gradients for segment $t$, the algorithm first retrieves the initial memory state $mem_{t}$ from the buffer. It then recomputes the forward pass for segment $t$ only, starting from $mem_{t}$ and using the input tokens $x_{t}$. This recomputation generates the necessary activations for calculating local gradients within segment $t$. The gradient flowing back from the subsequent segment $t+1$ (which is initialized using the stored $mem_{t+1}$) is then backpropagated through the recomputed segment $t$, including the update path for $mem_{t}$. This process is repeated backward for all segments. Algorithm 1 (Appendix D) provides a detailed pseudocode description. The “replay” aspect refers to this recomputation of the forward pass for each segment during the backward pass, using the stored astrocyte-inspired memory state as the starting point. AMRB offers significant memory efficiency and potential speed advantages. Unlike standard BPTT, which stores extensive activations, AMRB only caches the compact set of $M$ memory tokens passed between the $T_{seg}$ segments. Since $M$ is typically small, the memory footprint is drastically reduced. While AMRB involves recomputing activations during backpropagation, the associated memory saving often outweighs the recomputation cost for very long sequences, potentially leading to faster overall training (further details in Section 4) compared to standard BPTT.
@@
In summary, RMAAT introduces a novel astrocyte-inspired adaptive memory compression system. This is realized through the Memory Retention Factor derived from simulated astrocyte LTP, providing a principled, non-learned method for compressing and propagating context between segments. Our core hypothesis is that this principled compression enables a more efficient training strategy. By structuring the flow of information into a compressed set of memory tokens, we can forgo backpropagation through every token—the source of high memory costs in standard BPTT used in RMT—and instead use our highly memory-efficient AMRB algorithm. Our ablation study in Section 4.2 validates this critical synergy: removing the compression (a setup resembling RMT with memory replay) causes a significant accuracy drop. This demonstrates that our bio-inspired compression is crucial for making the memory-saving AMRB algorithm effective, and this combination is directly responsible for RMAAT’s gains in both efficiency and accuracy.
@@ 4.2 Ablation Studies
Memory Retention Factor (Contributions $1$ & $2$): Removing the retention factor significantly reduced accuracy on the Retrieval task (83.283.2% $\rightarrow$ 80.580.5%) without changing memory usage ($3.4$ GB), confirming its vital role in context compression derived from the LTP macro model.
@@
AMRB Training (Contribution $3$): Replacing AMRB with standard BPTT for the Retrieval task yielded similar accuracy but drastically increased peak memory usage ($\sim 4.4\times$, from $3.4$ GB to $15.0$ GB), demonstrating AMRB’s substantial memory efficiency benefits.
@@
Total Sequence Length: Evaluating performance on shorter total sequence lengths by reducing the number of segments (while keeping segment sequence length constant at $512$) resulted in significant accuracy drops in Retrieval task (e.g., 71.571.5% for $8$ segments [$4K$ total length], 65.365.3% for $4$ segments [$2K$ total length] vs. 83.283.2% for the baseline $16$ segments [$8K$ total length]), demonstrating that the model benefits from processing the full context length allowed by the segmentation strategy.
@@ 5 Conclusion
This work introduced the Recurrent Memory Augmented Astromorphic Transformer (RMAAT), demonstrating an effective approach to efficient long-sequence modeling by integrating computationally abstracted principles from astrocyte function. By incorporating astrocyte-inspired mechanisms for temporal memory compression and resource-efficient training (AMRB), RMAAT achieves state-of-the-art average accuracy on the diverse Long Range Arena benchmark. This strong performance, coupled with competitive memory efficiency compared to standard and recurrent baselines (Table 1), validates the potential of leveraging neuro-glial principles for challenging sequence tasks. While promising, the current evaluation is primarily focused on LRA; future work should explore broader domains, larger model scales, and deeper theoretical analysis comparing RMAAT to related sequence model formalisms. Investigating additional neuro-glial computational mechanisms (such as astrocyte-astrocyte communication) and developing specialized hardware implementations also present exciting avenues. In conclusion, RMAAT highlights the value of neuroscience-algorithm co-design, suggesting that astromorphic computing is a promising direction for developing powerful and efficient AI systems capable of handling complex, long-range sequential data.
@@ A.1 Neural Dynamics
$\tau_{n}\frac{dV_{i}(t)}{dt}=-\lambda(V_{i}(t)-V_{reset})+I_{i}(t)$ (S1)
@@ A.2 Synaptic Dynamics
$\tau_{s}\frac{ds_{ij}}{dt}=-\beta s_{ij}+\theta(x_{i})\theta(x_{j})+\psi(p_{ij}^{s})+c_{ij}$ (S3)
@@ A.3 Short-Term Astrocytic Process Dynamics (STP)
$\tau_{p}^{s}\frac{dp_{ij}^{s}}{dt}=-\gamma^{s}p_{ij}^{s}+\sum_{k,l=1}^{N}T_{ijkl}\psi(p_{kl}^{s})+d_{ij}$ (S4)
@@ A.4 Long-Term Astrocytic Process Dynamics (LTP)
$\tau_{p}^{l}\frac{dp_{ij}^{l}}{dt}=-\gamma^{l}p_{ij}^{l}+\kappa(s_{ij})$ (S5)
@@ A.5 Parameter Values for Simulation
The specific values used for the simulations presented in the main text (e.g., Figure 3) are listed in Table S1.
Table S1: Computational Neuroscience Model Hyperparameters Used in Simulations.
@@ B.2 Write Mode: Encoding Context
    $H_{neuron}=\sum_{t=1}^{N}\frac{1}{m}h_{t}^{T}v_{t}=\frac{1}{m}\phi(K)^{T}V\quad\in\mathbb{R}^{m\times d}$ (S6)
    This captures baseline Hebbian learning, linked to the $\theta(x_{i})\theta(x_{j})$ term (Eq. S3).
Astrocyte-Modulated Hebbian Weight ($H_{astro}$): This component models the astrocyte’s influence on the hidden-to-output connection, incorporating spatial context via a relative positional encoding matrix $R\in\mathbb{R}^{N\times m}$. The computation of $R$ itself, detailed in Section 3.2.3 and inspired by STP spatial dynamics ($T_{ijkl}$ in Eq. S4), involves transforming a base distance matrix ($r_{ij}=\exp(-\|\text{pos}_{i}-\text{pos}_{j}\|\times\text{scale})$) using learnable projections $M$ and $W_{rel}$ ($R=W_{rel}(MrM^{T})$).
@@
    $H_{astro}=\sum_{t=1}^{N}\frac{1}{m}\phi(r_{t})^{T}v_{t}=\frac{1}{m}\phi(R)^{T}V\quad\in\mathbb{R}^{m\times d}$ (S7)
    This functionally abstracts the astrocyte modulation term $\psi(p_{ij}^{s})$ (Eq. S3).
@@
    $g=\left(\sum_{t=1}^{N}\phi(k_{t})\right)^{\alpha}$ (S8)
@@ B.3 Read Mode: Retrieving Context
$L=\phi(Q)\left(\left(\frac{1}{m}(\phi(K)^{T}+\phi(R)^{T})V\right)\odot\left(\frac{1}{C}\right)\right)+X$ (S13)
@@ Appendix C $T_{ijkl}$ Formation and Visualization
Figure S1: Spatial layout of the $N=5$ neurons used in the simulation.
This appendix details the calculation and visualization of the spatial coupling tensor $T_{ijkl}$. This tensor is crucial in the Short-Term Astrocytic Process Dynamics (STP) described in Appendix A (Eq. S4), where it models the distance-dependent influence between different astrocyte processes associated with synapses $(i,j)$ and $(k,l)$. Understanding its structure helps motivate the bio-inspired relative positional encoding used in Section 3.2.3. The specific simulation results visualized in this appendix (e.g., $T_{ijkl}$ slices and $p^{s}_{ij}$ dynamics) were generated using a network size of $N=5$ and a neural bias parameter $b=0.1$, run for a duration of 50 seconds (representing one STP cycle), with all other model parameters set as detailed in Appendix A (Table S1).
Figure S2: Calculated midpoint positions for all possible synapses between the $N=5$ neurons.
@@ C.1 Distance Calculation
For simulation purposes, we first define the spatial layout of the neurons. For the example shown, we consider $N=5$ neurons arranged linearly in a 1D space, assigned coordinates for visualization (see Figure S1). A synapse $(i,j)$ connects postsynaptic neuron $i$ and presynaptic neuron $j$. We define the spatial position of synapse $(i,j)$ as the midpoint between the coordinates of neuron $i$ and neuron $j$. This results in a grid of $N\times N=25$ possible synapse locations for $N=5$ neurons (see Figure S2).
@@ C.2 Synapse Position Calculation
Next, we calculate the pairwise Euclidean distance, $\mathrm{distance}_{ij,kl}$, between the midpoint coordinates of every pair of synapses $(i,j)$ and $(k,l)$. This forms a distance matrix capturing the spatial separation between all potential synaptic interaction sites (see Figure S3).
Figure S3: Heatmap visualizing the pairwise Euclidean distances between all synapse midpoints for $N=5$, i.e., $\mathrm{distance}_{ij,kl}$.
@@ C.3 $T_{ijkl}$ Formula and Visualization
$T_{ijkl}=\exp(-\mathrm{distance}_{ij,kl}\times\mathrm{scale})$ (S14)
@@
Visualizing slices of the $T_{ijkl}$ tensor helps to understand the spatial interaction profile from a specific source synapse $(i,j)$ to all possible target synapses $(k,l)$. Figure S4 shows examples for source synapses located at the corner, edge, and center of the $5\times 5$ grid, for different values of $\mathrm{scale}$. Brighter colors indicate stronger influence (smaller distance or smaller $\mathrm{scale}$). Notice how the spatial extent of the influence changes significantly with the $\mathrm{scale}$ parameter.
@@ C.4 Impact on Astrocyte Dynamics and Link to Positional Encoding
Figure S4: Visualization of $T_{ijkl}$ slices showing interaction strength from different source synapse locations (rows: center, corner, edge) to all target synapses ($k$,$l$ grid) for varying ‘scale’ parameters (columns: $2.0,5.0,20.0$).
The formulation of $T_{ijkl}$ as an exponential decay of distance (Eq. S14) implies that closer synaptic processes have a stronger potential for direct influence. However, the ultimate astrocytic response, represented by the short-term process dynamics $p^{s}_{ij}$ (Eq. S4), is not solely determined by $T_{ijkl}$. It results from the complex interplay of neuronal activity ($x_{i},x_{j}$), synaptic facilitation ($s_{ij}$), and the integrated influence from all other astrocyte processes ($\sum_{k,l}T_{ijkl}\psi(p_{kl}^{s})$). Therefore, while $T_{ijkl}$ defines the strength of individual pairwise couplings, it is the simulation of the entire neuron-astrocyte network that reveals how these distance-dependent couplings translate into spatially modulated astrocytic responses over time.
As shown in Figure S5, these simulations demonstrate that synapses located centrally indeed tend to exhibit different temporal dynamics for $p^{s}_{ij}$ (e.g., higher peak and sustained activity) compared to those at corners or edges. This occurs because central locations benefit from a stronger integrated influence from a larger number of relatively closer neighbors, as dictated by the $T_{ijkl}$ coupling strengths. Changing the $\mathrm{scale}$ parameter in $T_{ijkl}$ further alters the range and strength of these interactions, consequently affecting the resulting $p^{s}_{ij}$ dynamics.
Figure S5: Simulated temporal dynamics of the short-term astrocyte process parameter $p^{s}_{ij}$ for synapses at different locations (center, corner, edge) under different spatial coupling ‘scale’ values ($2.0$, $5.0$, $20.0$). Simulation uses $N=5$, $b=0.1$.

### v2

@@ 3.4 AMRB Training Algorithm
Training recurrent architectures on long sequences via standard BPTT is often memory-prohibitive due to storing activations for the entire sequence length. To overcome this while leveraging RMAAT’s unique memory structure, we introduce the Astrocytic Memory Replay Backpropagation (AMRB) algorithm (Contribution $3$), an efficient training approach inspired by techniques for recurrent networks (Bellec et al., 2019; Meng et al., 2023)The core idea of AMRB is to avoid storing all intermediate activations within each segment during the forward pass. Instead, it leverages the persistent, compressed Memory Tokens ($mem_{t}$) described in Section 3.3. During the forward pass through $T_{seg}$ segments, only the sequence of memory token states ($mem_{1},mem_{2},...,mem_{T_{seg}+1}$) passed between segments is stored in a replay buffer. During the backward pass, gradients are computed segment by segment. To calculate the gradients for segment $t$, the algorithm first retrieves the initial memory state $mem_{t}$ from the buffer. It then recomputes the forward pass for segment $t$ only, starting from $mem_{t}$ and using the input tokens $x_{t}$. This recomputation generates the necessary activations for calculating local gradients within segment $t$. The gradient flowing back from the subsequent segment $t+1$ (which is initialized using the stored $mem_{t+1}$) is then backpropagated through the recomputed segment $t$, including the update path for $mem_{t}$. This process is repeated backward for all segments. Algorithm 1 below provides a detailed pseudocode description. The “replay” aspect refers to this recomputation of the forward pass for each segment during the backward pass, using the stored astrocyte-inspired memory state as the starting point.
Input: $m_{1}$: Initial memory state (input to step $t=1$).
Output: $m_{T+1}$: Updated memory state (output after step $t=T$).
1
// Store initial input state $m_{1}$
2
/* Forward Pass */
3 for $t=1$ to $T$ do
  $m^{\prime}_{t+1}\leftarrow\text{Model}(x_{t},m_{t})$ ;
  // Compute intermediate state (no grad)
  // Apply retention factor
4  if $t<T$ then
     Append $m_{t+1}$ to replay_bufferreplay\_buffer ;
     // Store input state $m_{t+1}$ for next step’s recomputation
5
6
/* Backward Pass */
Initialize $\nabla m_{T+1}\leftarrow 0$ ;
// Init gradient for state after last step
7 for $t=T$ to 11 do
  Retrieve $m_{t}$ from replay_bufferreplay\_buffer ;
  // Get input state for segment $t$
  $m^{\prime}_{t+1},o_{t}\leftarrow\text{Model}(x_{t},m_{t})$ ;
  // Recompute segment $t$ (track grads)
  Compute loss $L_{t}\leftarrow\text{loss\_function}(o_{t})$ ;
  // Compute loss for current step
  Perform backpropagation: $L_{t}.\text{backward()}$ ;
  // Compute param grads $\partial L_{t}/\partial\theta_{t}$, etc.
  // Compute $\nabla m_{t}$ via chain rule
8
9Save $m_{T+1}$ for the next rollout’s update
Algorithm 1 Astrocytic Memory Replay Backpropagation (AMRB)
AMRB offers significant memory efficiency and potential speed advantages. Unlike standard BPTT, which stores extensive activations, AMRB only caches the compact set of $M$ memory tokens passed between the $T_{seg}$ segments. Since $M$ is typically small, the memory footprint is drastically reduced. While AMRB involves recomputing activations during backpropagation, the associated memory saving often outweighs the recomputation cost for very long sequences, potentially leading to faster overall training (further details in Section 4) compared to standard BPTT.
@@
In summary, RMAAT introduces a novel astrocyte-inspired adaptive memory compression system. This is realized through the Memory Retention Factor derived from simulated astrocyte LTP, providing a principled, non-learned method for compressing and propagating context between segments. Our core hypothesis is that this principled compression enables a more efficient training strategy. By structuring the flow of information into a compressed set of memory tokens, we can forgo backpropagation through every token—the source of high memory costs in standard BPTT used in RMT—and instead use our highly memory-efficient AMRB algorithm. Our ablation study in Section 4.2 validates this critical synergy: removing the compression causes a significant accuracy drop. This demonstrates that our bio-inspired compression is crucial for making the memory-saving AMRB algorithm effective, and this combination is directly responsible for RMAAT’s gains in both efficiency and accuracy.
@@ 4.2 Ablation Studies
Memory Retention Factor (Contributions $1$ & $2$): Removing the retention factor significantly reduced accuracy on the Retrieval task (83.283.2% $\rightarrow$ 80.580.5%) without changing memory usage ($3.4$ GB), confirming its vital role in context compression derived from the LTP macro model. This effect is consistent across modalities: on the Text ($4K$) task, accuracy decreases from 65.965.9% to 64.964.9% when the retention factor is removed.
AMRB Training (Contribution $3$): Replacing AMRB with standard BPTT yielded similar accuracy but drastically increased peak memory usage on both Retrieval ($3.4$ GB $\rightarrow$ $15.0$ GB, $\sim 4.4\times$) and Text ($5.1$ GB $\rightarrow$ $22.0$ GB, $\sim 4.3\times$), demonstrating AMRB’s memory efficiency benefits.
RLT + AMRB Ablation (Astromorphic Components): We also evaluate an RLT + AMRB variant, which applies the Memory Retention Factor and AMRB to the Recurrent Linear Transformer (RLT) baseline (standard linearized attention, without $H_{astro}$ or $P$). On Retrieval ($8K$), RLT + AMRB attains 79.279.2% accuracy with peak memory $\sim 3.4$ GB and RMAAT-like throughput, versus 78.478.4% and $12.1$ GB for RLT and 83.283.2% and $3.4$ GB for RMAAT.
@@
Total Sequence Length: Evaluating performance on shorter total sequence lengths by reducing the number of segments (while keeping segment sequence length constant at $512$) resulted in significant accuracy drops in the Retrieval task (e.g., 71.571.5% for $8$ segments [$4K$ total length], 65.365.3% for $4$ segments [$2K$ total length] vs. 83.283.2% for the baseline $16$ segments [$8K$ total length]), demonstrating that the model benefits from processing the full context length allowed by the segmentation strategy.
@@ 5 Conclusion
This work introduced the Recurrent Memory Augmented Astromorphic Transformer (RMAAT), demonstrating an effective approach to efficient long-sequence modeling by integrating computationally abstracted principles from astrocyte function. By incorporating astrocyte-inspired mechanisms for temporal memory compression and resource-efficient training (AMRB), RMAAT achieves strong average accuracy on the diverse Long Range Arena benchmark. This performance, coupled with competitive memory efficiency compared to standard and recurrent baselines (Table 1), validates the potential of leveraging neuro-glial principles for challenging sequence tasks within this benchmark setting. While promising, the current evaluation is primarily focused on LRA; future work should explore broader domains (including richer vision and multimodal settings), larger model scales, and extensions to true streaming or online processing where the total sequence length is not known in advance, as well as deeper theoretical analysis comparing RMAAT to related sequence model formalisms. Investigating additional neuro-glial computational mechanisms (such as astrocyte-astrocyte communication) and developing specialized hardware implementations also present exciting avenues. In conclusion, RMAAT highlights the value of neuroscience-algorithm co-design, suggesting that astromorphic computing is a promising direction for developing powerful and efficient AI systems capable of handling complex, long-range sequential data.
Acknowledgments
Research was sponsored in part by the Army Research Office and was accomplished under Grant Number W911NF-24-1-0127. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was also supported partially by the National Science Foundation under award No. EFRI BRAID #2318101.
@@ A.1 Neural Dynamics
$\tau_{n}\frac{dV_{i}(t)}{dt}=-\lambda(V_{i}(t)-V_{reset})+I_{i}(t)$ (7)
@@ A.2 Synaptic Dynamics
$\tau_{s}\frac{ds_{ij}}{dt}=-\beta s_{ij}+\theta(x_{i})\theta(x_{j})+\psi(p_{ij}^{s})+c_{ij}$ (9)
@@ A.3 Short-Term Astrocytic Process Dynamics (STP)
$\tau_{p}^{s}\frac{dp_{ij}^{s}}{dt}=-\gamma^{s}p_{ij}^{s}+\sum_{k,l=1}^{N}T_{ijkl}\psi(p_{kl}^{s})+d_{ij}$ (10)
@@ A.4 Long-Term Astrocytic Process Dynamics (LTP)
$\tau_{p}^{l}\frac{dp_{ij}^{l}}{dt}=-\gamma^{l}p_{ij}^{l}+\kappa(s_{ij})$ (11)
@@ A.5 Parameter Values for Simulation
The specific values used for the simulations presented in the main text (e.g., Figure 3) are listed in Table 3.
Table 3: Computational Neuroscience Model Hyperparameters Used in Simulations.
@@ B.2 Write Mode: Encoding Context
    $H_{neuron}=\sum_{t=1}^{N}\frac{1}{m}h_{t}^{T}v_{t}=\frac{1}{m}\phi(K)^{T}V\quad\in\mathbb{R}^{m\times d}$ (12)
    This captures baseline Hebbian learning, linked to the $\theta(x_{i})\theta(x_{j})$ term (Eq. 9).
Astrocyte-Modulated Hebbian Weight ($H_{astro}$): This component models the astrocyte’s influence on the hidden-to-output connection, incorporating spatial context via a relative positional encoding matrix $R\in\mathbb{R}^{N\times m}$. The computation of $R$ itself, detailed in Section 3.2.3 and inspired by STP spatial dynamics ($T_{ijkl}$ in Eq. 10), involves transforming a base distance matrix ($r_{ij}=\exp(-\|\text{pos}_{i}-\text{pos}_{j}\|\times\text{scale})$) using learnable projections $M$ and $W_{rel}$ ($R=W_{rel}(MrM^{T})$).
@@
    $H_{astro}=\sum_{t=1}^{N}\frac{1}{m}\phi(r_{t})^{T}v_{t}=\frac{1}{m}\phi(R)^{T}V\quad\in\mathbb{R}^{m\times d}$ (13)
    This functionally abstracts the astrocyte modulation term $\psi(p_{ij}^{s})$ (Eq. 9).
@@
    $g=\left(\sum_{t=1}^{N}\phi(k_{t})\right)^{\alpha}$ (14)
@@ B.3 Read Mode: Retrieving Context
$L=\phi(Q)\left(\left(\frac{1}{m}(\phi(K)^{T}+\phi(R)^{T})V\right)\odot\left(\frac{1}{C}\right)\right)+X$ (19)
@@ Appendix C $T_{ijkl}$ Formation and Visualization
Figure 5: Spatial layout of the $N=5$ neurons used in the simulation.
This appendix details the calculation and visualization of the spatial coupling tensor $T_{ijkl}$. This tensor is crucial in the Short-Term Astrocytic Process Dynamics (STP) described in Appendix A (Eq. 10), where it models the distance-dependent influence between different astrocyte processes associated with synapses $(i,j)$ and $(k,l)$. Understanding its structure helps motivate the bio-inspired relative positional encoding used in Section 3.2.3. The specific simulation results visualized in this appendix (e.g., $T_{ijkl}$ slices and $p^{s}_{ij}$ dynamics) were generated using a network size of $N=5$ and a neural bias parameter $b=0.1$, run for a duration of 50 seconds (representing one STP cycle), with all other model parameters set as detailed in Appendix A (Table 3).
Figure 6: Calculated midpoint positions for all possible synapses between the $N=5$ neurons.
@@ C.1 Distance Calculation
For simulation purposes, we first define the spatial layout of the neurons. For the example shown, we consider $N=5$ neurons arranged linearly in a 1D space, assigned coordinates for visualization (see Figure 5). A synapse $(i,j)$ connects postsynaptic neuron $i$ and presynaptic neuron $j$. We define the spatial position of synapse $(i,j)$ as the midpoint between the coordinates of neuron $i$ and neuron $j$. This results in a grid of $N\times N=25$ possible synapse locations for $N=5$ neurons (see Figure 6).
@@ C.2 Synapse Position Calculation
Next, we calculate the pairwise Euclidean distance, $\mathrm{distance}_{ij,kl}$, between the midpoint coordinates of every pair of synapses $(i,j)$ and $(k,l)$. This forms a distance matrix capturing the spatial separation between all potential synaptic interaction sites (see Figure 7).
Figure 7: Heatmap visualizing the pairwise Euclidean distances between all synapse midpoints for $N=5$, i.e., $\mathrm{distance}_{ij,kl}$.
@@ C.3 $T_{ijkl}$ Formula and Visualization
$T_{ijkl}=\exp(-\mathrm{distance}_{ij,kl}\times\mathrm{scale})$ (20)
@@
Visualizing slices of the $T_{ijkl}$ tensor helps to understand the spatial interaction profile from a specific source synapse $(i,j)$ to all possible target synapses $(k,l)$. Figure 8 shows examples for source synapses located at the corner, edge, and center of the $5\times 5$ grid, for different values of $\mathrm{scale}$. Brighter colors indicate stronger influence (smaller distance or smaller $\mathrm{scale}$). Notice how the spatial extent of the influence changes significantly with the $\mathrm{scale}$ parameter.
@@ C.4 Impact on Astrocyte Dynamics and Link to Positional Encoding
Figure 8: Visualization of $T_{ijkl}$ slices showing interaction strength from different source synapse locations (rows: center, corner, edge) to all target synapses ($k$,$l$ grid) for varying ‘scale’ parameters (columns: $2.0,5.0,20.0$).
The formulation of $T_{ijkl}$ as an exponential decay of distance (Eq. 20) implies that closer synaptic processes have a stronger potential for direct influence. However, the ultimate astrocytic response, represented by the short-term process dynamics $p^{s}_{ij}$ (Eq. 10), is not solely determined by $T_{ijkl}$. It results from the complex interplay of neuronal activity ($x_{i},x_{j}$), synaptic facilitation ($s_{ij}$), and the integrated influence from all other astrocyte processes ($\sum_{k,l}T_{ijkl}\psi(p_{kl}^{s})$). Therefore, while $T_{ijkl}$ defines the strength of individual pairwise couplings, it is the simulation of the entire neuron-astrocyte network that reveals how these distance-dependent couplings translate into spatially modulated astrocytic responses over time.
As shown in Figure 9, these simulations demonstrate that synapses located centrally indeed tend to exhibit different temporal dynamics for $p^{s}_{ij}$ (e.g., higher peak and sustained activity) compared to those at corners or edges. This occurs because central locations benefit from a stronger integrated influence from a larger number of relatively closer neighbors, as dictated by the $T_{ijkl}$ coupling strengths. Changing the $\mathrm{scale}$ parameter in $T_{ijkl}$ further alters the range and strength of these interactions, consequently affecting the resulting $p^{s}_{ij}$ dynamics.
Figure 9: Simulated temporal dynamics of the short-term astrocyte process parameter $p^{s}_{ij}$ for synapses at different locations (center, corner, edge) under different spatial coupling ‘scale’ values ($2.0$, $5.0$, $20.0$). Simulation uses $N=5$, $b=0.1$.